The meeting commenced at 1037 hours with the following in attendance.
|Mick Day (Chairman)||-|
|Boyd Potts (Secretary)||Orpheus C.C.|
|Ben Lovett (Conservation Officer)||Morgannwg C.C.|
|Ali Garman (Survey Officer)||Draenan Diggers|
|Chris Densham (Trustee)||Oxford University C.C.|
|John Cooper||Chelsea S.S.|
|Martin Laverty||Grwp Ogofeydd Garimpeiros|
|Ian Wilton-Jones||Brynmawr C.C.|
|Charles Bailey (Trustee)||-|
|Also present as an observer was Jon Jones (Draenan Diggers).|
Confirmed the meeting was quorate and opened proceedings.
Elsie Little (CCC), Spencer Drew (Treasurer), Andy Farrant (Geological Recorder), Rhian Hicks (Biological Recorder), Sue Mabbett (Access Officer), Owen Clarke (Cwmbran C.C.), Pete Bolt (BEC).
3(a)It was agreed unanimously, 'That the minutes be accepted as a true record.'
Item 4(b). No request yet received for payment of legal fees.
Item 4(c). Hereford CC's status as a key holder not recorded, as SM not present.
Item 4(d). Taping notice still to be placed.
Item 8(d). E.Mail circulation working well. The Minutes are now on the website.
(i). Circulation of correspondence between the committee, JS/AM and JC was recorded. Discussion was held on future over the matter. Suggested by AG, and agreed, that we forward all the data to JS and ask for comments before we place all our information on the website.
(ii). Noted that Conservation Policy and other items were passed to Mike Simms for a meeting in Ireland also attended by BL. BL suggested we sort out the Conservation Policy and then publish on the website. The Policy to be ratified at the next meeting.
(iii). Still no news about the Rolex award. It was pointed out that the Referees were not aware of the Group's existence when the application was made. AF has stated he will keep the Group informed of any progress.
(iv). Correspondence and discussion noted about the possible World Heritage Site and its relevance to the cave system.
Not present but had forwarded Accounts which were circulated at the meeting. Balance in account at 31st May 2000 was £906.48. All agreed, 'That the Treasurer's report be accepted.' Discussion on the 'Future Cash Flow Projection.' It was agreed that we keep this under review and also look for access to other sources of funding for the future.
Not present but report was circulated at the meeting. Request that the Book Box fitting at the Lamb and Fox be expedited. It was agreed that after discussion with Brian (the Landlord), BL and JJ will take the necessary action. The Book details to be arranged by SM.
(i) Diesel leak in February. Traced to a lorry running off the road. Welsh Water informed and appreciative of the information. Possible that a groundwater protection zone may be set up round the cave drainage area; this would expedite pollution control and help minimise surface pollution. There was a long discussion on dumping and pollution. Noted that dye tracing should be done with Welsh Water approval. We need to keep a good working relationship but try to avoid possible restrictions.
(ii) Taping. This is ongoing. Tape stocks in the cave will be kept supplied and more tape is available when needed. Clubs willing to tape are asked to contact BL over areas needing tape. Comment from IW-J that stals still being touched where there is no necessity.
(iii) Two reports of Second Entrance turf being removed. At the end of March and again in the middle of May. On the second occasion the persons were seen and asked by Brian for their reasons. They claimed (falsely) to have been given permission. Discussion held on why this was being done. BP to approach PD Conservation Ltd for their comments and to discuss what action we might take against the perpetrators.
(iv) BL to take action over the Taping notice agreed at the last meeting.
No new data received directly since the last meeting. All requests for use of data have been dealt with. The total cave length held as G2 data is 48,565m plus 386m not compiled. Other areas have been discovered and surveyed by the G5 surveyors but no data supplied to the Group. As said in the Secretary's report; AG is making all G2 data available to the G5 surveyors for their comments before it goes on the website. Comments made that it was a pity that the G5 surveyors were not working within the Group rather than on the outside.
Not present but has forwarded comments.
(i) Request from an external source to use the biological data from the cave. This was agreed by the Group.
(ii) Request for more information for the website, it still lacks content. Comment that the website needs to be available as widely as possible.
(iii) Committee agreed that E.Mail details should go on the website, but not addresses and phone numbers. The exception to this may be the Access Officer but that needs to be checked with SM.
Not in attendance but had commented that there was no new information at present. Comments from OC noted over his work on sites linked to the cave system.
Access to Second Entrance was discussed and BP stated that he was keeping the Landowners informed of any possible requests for access. The present policy would continue unless a change was agreed between PD Conservation and the Group.
7(a) Mine entrance near Keepers Pond noted. Outside our remit but suggested that TG and the Hang Gliders Group be informed.
7(b) Discussion over the route down to the present entrance. Agreed it seems stable at present but needs careful monitoring for possible future maintenance.
8(a) The ladder on the Balcony Pitch needs replacement or major repairs. Request that the next group into this area place a warning notice and advising use of a lifeline. Agreed we publicise this in Descent, Speleoscene, etc. AG to action. BL to organise disposal or repair of ladder as soon as possible.
8(b) After discussion the following was Proposed by ML and Seconded by BP; 'The Committee recognises that cavers may install aids in the cave system but takes no responsibility for their maintenance or removal. Cavers use any aids at their own risk.' This was passed unanimously.
8(c) Discussion was held on what aids are already in the cave and whether any lists, official or unofficial, should be made. Agreed that no lists should be produced by this Group.
There was none.
This was arranged and confirmed as Sunday 26th November 2000. To be held at the Gwesty Bach, Brynmawr, at 10.30am.
The meeting closed at 1330 hours.
Note: Treasurer's and Access Officer's Reports can be supplied by post if requested.